# RL3: Boosting Meta Reinforcement Learning via RL inside RL2

### v5

@@ 4.1 Theoretical Justification
Here, we consider the interpretation of meta-RL as performing RL on a partially observable Markov decision process (POMDP) in which the partially observable state factor is the identity of the object-level MDP. Without loss of generality, all analysis assumes the infinite horizon setting. We will denote meta-level entities, belonging in this case to a POMDP, with an overbar. For example, we have a meta-level value function $\bar{V}$ and a meta-level belief $\bar{b}$.
First, we show a basic result, that the optimal meta-level value function is upper bounded by the object-level Q-value estimates in the limit.
Proof: Given a task distribution $\mathcal{M}$, then for state $s$, there exists a maximum object-level optimal value function $V^{*}_{max}(s)$, corresponding to some MDP $M_{max}\in\mathcal{M}$, such that for all MDPs $M_{i}\in\mathcal{M}$, $V^{*}_{max}(s)\geq V^{*}_{i}(s)$. The expected cumulative discounted reward experienced by the agent cannot be greater than the most optimistic value function over all tasks, since $\bar{V}^{*}(\bar{b})$ is a weighted average of individual value functions $V^{\pi_{\theta}}(s)$, which are themselves upper bounded by $V^{*}_{max}(s)$. Thus,
@@
Next, we see that combining the asymptotic accuracy of Q-estimates and Equation (2) yields
$\lim_{t\rightarrow\infty}\max_{a\in A,M_{i}\in\mathcal{M}}Q_{i}^{t}(s,a)\geq\bar{V}^{*}(\bar{b})\hskip 14.22636pt\forall s\in S.\qed$ (3)
Furthermore, it follows if the meta-level observation $\bar{\omega}$ includes Q-value estimates of the current task $M_{i}$, it can be shown that as $t\rightarrow\infty$, the optimal meta-value function approaches the optimal value function for the current task, i.e., for any $\epsilon>0$, there exists $\kappa\in\mathbb{N}$ such that for $t\geq\kappa$,
$\Big{|}\max_{a\in A}\Big{[}Q_{i}^{t}(s,a)\Big{]}-\bar{V}^{*}(\bar{b})\Big{|}\leq\epsilon\hskip 14.22636pt\forall s\in S.$ (4)
Equation 4 (proof in Appendix A.3) shows that for $t\geq\kappa$, acting greedily w.r.t. $Q_{i}^{*}$ leads to Bayes-optimal behavior, and knowing the Bayes-optimal belief over tasks is not required, implying that the experience history can be ignored at that point. Moreover, it follows from equation 4 that for $t<\kappa$,
where error $\varepsilon_{i}(\Upsilon)$ is the error in Q-value estimates. While this error will diminish as $t\rightarrow\infty$, in the short run, a function $f(\Upsilon)$ could be learned to either estimate the error or estimate $\bar{V}^{*}(\bar{b})$ entirely.
The better performance of RL3 could be explained by either error $\varepsilon_{i}(\Upsilon)$ being simpler to estimate, or, the meta-agent behavior being more robust to errors in estimates of $\varepsilon_{i}(\Upsilon)$ when Q-estimates are supplied directly as inputs, than to errors in a more complicated approximation of $\bar{V}^{*}(\bar{b})$. Moreover, this composition benefits from the fact that the convergence rate for Q-estimates suggests a natural, predictable rate of shifting reliance from $f(\Upsilon)$ to $Q^{t}_{i}(s)$ as $t\rightarrow\infty$. However, we do not bake this structure into the network and instead let it implicitly learn how much to use the Q-estimates.
Finally, we note that near-perfect function approximation of $\bar{V}^{*}(\bar{b})$ as $t\rightarrow\infty$ reduces error in meta-value function approximation for all preceding belief states, as meta-values for consecutive belief states $\bar{b}$ and $\bar{b}^{\prime}$ are linked through the Bellman equation for BAMDPs (see details in Appendix A.3)
$\bar{V}^{*}(\bar{b})=\max_{a\in A}\Big{[}\!\sum_{M_{i}\in\mathcal{M}}\!\!\bar{b}(i)R_{i}(s,a)+\gamma\!\sum_{\bar{\omega}\in\bar{\Omega}}\bar{O}(\bar{\omega}|\bar{b},a)\bar{V}^{*}(\bar{b}^{\prime})\Big{]}.$ (6)
This dependency helps meta-training in RL3 with temporal-difference based learning algorithms. Without conditioning on Q-estimates, error in $\bar{V}^{*}(\bar{b})$ would instead increase as $t\rightarrow\infty$, as the meta-critic would be conditioned on a larger history, which could destabilize the meta-value learning for all preceding belief states during meta-training.
4.2 Implementation
Implementing RL3 involves simply replacing each MDP in the task distribution with a corresponding value-augmented MDP (VAMDP) and solving the resulting VAMDP distribution using RL2. Each VAMDP has the same action space and reward function as the corresponding MDP. The value augmented state $\hat{s}_{t}\in S\times\mathbb{R}^{k}\times\mathbb{I}^{k}$ includes the object level state $s_{t}$, $k$ real values and $k$ integer values for the Q-estimates ($Q^{t}(s_{t},a)$) and action counts ($N^{t}(s_{t},a)$) for each of the $k$ actions.In practice, we provide action advantages along with the max Q-value (value function) instead of Q-estimates. When the object-level state space $S$ is discrete, $s_{t}$ needs to be represented as an $|S|$-dimensional one-hot vector. Note that the value augmented state space is continuous. In the VAMDP transition function, the object-level state $s$ has the same transition dynamics as the original MDP, while the dynamics of Q-estimates are a function of $T$, $R$, and the specific object-level RL algorithm used for estimating Q-values. An episode of the VAMDP spans the entire interaction period with the corresponding MDP, which may include multiple episodes of the MDP, as Q-estimates continue to evolve beyond episode boundaries. In code, a VAMDP RL environment is implemented as a wrapper over a given MDP environment. The pseudocode, additional implementation details and hyperparameters for RL2 and RL3 are mentioned in Appendix
@@ 5 Experiments
We compare RL3 to our enhanced implementation of RL2. In our implementation, we replace LSTMs with transformers in both the meta-actor and meta-critic for the purpose of mapping experiences to actions and meta-values, respectively. This is done to improve RL2 ’s ability to handle long-term dependencies instead of suffering from vanishing gradients. Moreover, RL2-transformer trains significantly faster than RL2-LSTM. Second, we include in the state space the total number of interaction steps and the total number of steps within each episode during a meta-episode (see Fig. 1). Third, we use PPO [2] for training the meta actor-critic, instead of TRPO [54]. These modifications and other minor-implementation details incorporate the recommendations made by Ni et al. [6], who show that model-free recurrent RL is competitive with other state-of-the-art meta RL approaches such as VariBAD [7], if implemented properly. RL3 simply applies the modified version of RL2 to the distribution of value-augmented MDPs explained in section 4.2. Within each VAMDP, our choice of object-level RL is a model-based algorithm to maximize data efficiency – we estimate a tabular model of the environment and run finite-horizon value-iteration using the model. Once again, we emphasize that the core of our approach, which is augmenting MDP states with action-value estimates, is not inherently tied to RL2 and is orthogonal to most other meta-RL research. VAMDPs can be plugged into any base meta-RL algorithm with a reasonable expectation of improving it.
@@
Bernoulli Bandits: We use the same setup described by Duan et al. [4] with $k=5$ arms. To test OOD generalization, we generate bandit tasks by sampling success probabilities from $\mathcal{N}(0.5,0.5)$. We should note that this is an easy domain and serves as a sanity check to ensure that Q-value estimates do not hurt RL3, causing inferior performance.
Random MDPs: We use the same setup described by Duan et al. [4]. The MDPs have 10 states, 5 actions, and task horizon 10. The rewards and transition probabilities are drawn from a normal and a flat Dirichlet distribution ($\alpha=1.0$), respectively. OOD test MDPs use Dirichlet $\alpha=0.25$. We should note that this domain is particularly challenging for RL3 due to the high degree of stochasticity and thus the slower convergence rate of Q-estimates.
@@ 6 Results
In summary, we observe that beyond matching or exceeding the performance of RL2 in all test domains i) RL3 shows better OOD generalization, which we attribute to the increased generality of the Q-value representation, ii) the advantages of RL3 increase with longer interactions periods and less stochastic tasks, which we attribute to the increased accuracy of the Q-value estimates, iii) RL3 performs well even with coarse-grained object-level RL over abstract states with substantial computational savings, showing minimal drop in performance in most cases, and iv) RL3 shows faster meta-training.
Figure 3: Results for the MDPs and GridWorlds domains. Figure 3a shows the average cumulative reward (negligible standard error) earned as a fraction of the oracle policy for in-distribution (solid) and OOD (dashed) tasks; Figure 3b shows the fraction of RL2-transformer meta-training iterations that RL3 requires (variance is insignificant across seeds) to match RL2-transformer performance or fully converge, both as functions of the adaptation period. Note the log horizontal axis on both plots. Figure 3c shows the average cumulative reward ($\pm$ standard error) earned by RL2, RL3 , and RL3-coarse agents on several variations of the GridWorlds domain.
@@
MDPs: Figures 3a and 3b show the results for the MDPs domain. In Figure 3a, we see that for relatively short budgets, $H\leq 500$, both RL3 and RL2-transformer perform comparably on in-distribution problems, with RL3 performing slightly better on OOD tasks. We suspect that, due to the short budgets and highly stochastic domain, Q-estimates do not converge enough to be very useful for RL3 . However, as the budget increases, we see that RL3 continues to improve while RL2-transformer actually becomes worse and the performance gap on both in-distribution and OOD tasks becomes significant. Overall, we see that RL3 preserves asymptotic scaling properties of traditional RL while simultaneously maintaining strong OOD performance. Moreover RL3 it is able to learn meta-policies much more efficiently. Figure 3b shows the number of iterations of PPO RL3 takes to converge completely, as well as to match the performance of RL2-transformer, measured as a fraction of the time it takes for RL2-transformer to converge. This advantage of RL3 is again most pronounced for longer adaptation periods, but we still do observe significant meta-training speedup on even moderate ones. Overall, it is clear that as adaptation periods grow, RL3 achieves nearer-to-optimal policies in a fraction of the meta-training time and maintains better OOD generalization.
GridWorlds: Fig 3c shows the results for the GridWorld domain. On 11x11 grids with $H=250$, RL3 significantly outperforms RL2. On 13x13 grids with $H=350$, the performance margin is even greater, showing that while RL2-transformer struggles with a greater number of states, a longer adaptation period and more long-term dependencies, RL3 can take advantage of the Q-estimates to overcome the challenge. We also test the OOD generalization of both approaches in different ways by varying certain parameters of the 13x13 grids, namely, increasing the obstacle density (Dense), making actions on non-water tiles deterministic (Deterministic), increasing the number of wet ‘W’ tiles (Watery), increasing the number of danger ‘X’ tiles (Dangerous) and having the goal only in the corners (Corner). On all variations, RL3 continues to significantly outperform RL2. In a particularly interesting outcome, both approaches show improved performance on the Deterministic variation. However, RL3 gains 80% more points than RL2 , which is likely because Q-estimates converge faster on this less stochastic MDP and therefore provide greater help to RL3 . Conversely, in the Watery variation, which is more stochastic, both RL2 and RL3 lose roughly equal number of points. Overall, in each case, RL3-coarse significantly outperforms RL2-transformer. In fact, it performs on par with RL3, even outperforming it on Corner variation, except on the canonical 13x13 case and its Deterministic variation, where it scores about 90% of the scores for RL3. Finally, we see similar meta-training speedups where RL3 requires just $50\%$ and $30\%$ of the total iterations to match the performance of RL2-transformer on the 11x11 and 13x13 grids, respectively.
@@
Fig. 4 shows a sequence of snapshots of a meta-episode where the trained RL3 agent is interacting with an instance of a 13x13 grid. The first snapshot shows the agent just before reaching the goal for the first time. Prior to the first snapshot, the agent had explored many locations in the grid. The second snapshot shows the next episode just after the agent finds the goal, resulting in value estimates being updated using object-level RL for all visited states. Snapshot 3 shows the agent consequently using the Q-estimates to navigate to the goal presumably by choosing high-value actions. The agent also explores several new nearby states for which it does not have Q-estimates. Snapshot 4 shows the final Q-value estimates. A set of short videos of the GridWorld environment, showing both RL2 and RL3 agents solving the same set of problem instances, can be found at https://youtu.be/eLA_S1BQUYM.
Computation Overhead Considerations: As mentioned earlier, for implementing object-level RL, we use model estimation followed by finite-horizon value-iteration to obtain Q-estimates. The computation overhead is negligible for Bandits (5 actions, task horizon = 1) and very little for the MDPs domain (10 states, 5 actions, task horizon 10). For 13x13 GridWorlds (up to 169 states, 5 actions, task horizon = 350), RL3 takes approximately twice the computation time of RL2 per meta-episode. However, RL3-coarse requires only 10% overhead while still outperforming RL2 and retaining more than 90% of the performance of RL3. This demonstrates the utility of state abstractions in RL3 for scaling. Finally, the meta-training sample efficiency demonstrated by RL3 translates directly to wall-time efficiency as training is dominated by gradient computation, not value iteration during data collection in PPO. Our implementation is available at https://github.com/bhatiaabhinav/RL3.
@@ 7 Limitations and Conclusion
To conclude, in this paper, we introduced RL3, a principled hybrid approach that combines the strengths of traditional RL and meta-RL and provides a more robust and efficient meta-RL algorithm. We advanced intuitive and theoretical arguments regarding its suitability for meta-RL and presented empirical evidence to validate those ideas. Specifically, we demonstrated that RL3 holds potential to enhance long-term performance, generalization on out-of-distribution tasks and reducing meta-training time. In future work, we plan to explore extending RL3 to handle continuous state spaces.
@@ Acknowledgments
A Proofs
@@ A.1 Bayes Optimality of Q-value Estimates in Bernoulli Multi-armed Bandits
Let the number of times arm $k$ is pulled up to time $T$ be $N_{ik}^{T}$, and the number of successes associated with pulling arm $k$ up to time $T$ be $q_{ik}^{T}$. Given that this is an MDP with just a single state and task horizon of 1, the $Q$-estimate associated with arm $k$ is just the average reward for that action, which is the ratio of successes to counts associated with that action i.e., $Q_{ik}^{T}=\frac{q_{ik}^{T}}{N_{ik}^{T}}$. To reduce the clutter in the notation, we will drop the superscript $T$ for the rest of the subsection.
Now,

### v6

@@ 4.1 Theoretical Justification
Throughout this section, without loss of generality, we assume: i) an infinite-horizon setting, ii) that $\mathcal{M}$ denotes the support set of possible MDPs over which a distribution is defined, iii) all MDPs share the same state and action spaces, and iv) rewards are uniformly bounded by $|R_{i}(s,a)|\leq R_{\max}$. We consider the interpretation of Meta-RL as performing RL on a partially observable Markov decision process (POMDP) in which the partially observable factor is the identity of the object-level MDP. We denote meta-level entities (i.e., those belonging to the POMDP) with an overbar; for example, the meta-level value function $\bar{V}$ and the meta-level belief $\bar{b}$. First, we establish that the optimal meta-level value function is upper bounded by the object-level $Q$-values.
Proof. For a given state $s$, there exists a maximum object-level optimal value function $V^{*}_{\max}(s)$ corresponding to some MDP $M_{\max}\in\mathcal{M}$, such that for all MDPs $M_{i}\in\mathcal{M}$, $V^{*}_{\max}(s)\geq V^{*}_{i}(s)$. The expected cumulative discounted reward experienced by the agent cannot exceed this most optimistic value, since $\bar{V}^{*}(\bar{b})$ is a weighted average of the individual value functions $V^{\pi_{\theta}}(s)$, each of which is upper bounded by $V^{*}_{\max}(s)$. Thus,
$\max_{M_{i}\in\mathcal{M}}V^{*}_{i}(s)\ \geq\ \bar{V}^{*}(\bar{b})\quad\forall s\in S.$ (2)
Combining the asymptotic accuracy of $Q$-estimates with Equation 2 yields
$\lim_{t\to\infty}\max_{a\in A,M_{i}\in\mathcal{M}}Q_{i}^{t}(s,a)\geq\bar{V}^{*}(\bar{b})\quad\forall s\in S.\qed$ (3)
Furthermore, if the meta-level observation $\bar{\omega}$ includes the $Q$-value estimates of the current task $M_{i}$, it can be shown that as $t\to\infty$, the optimal meta-value function approaches the optimal value function for the current task. Specifically, for any $\epsilon>0$, there exists $\kappa\in\mathbb{N}$ such that for all $t\geq\kappa$,
$\Big{|}\max_{a\in A}\Big{[}Q_{i}^{t}(s,a)\Big{]}-\bar{V}^{*}(\bar{b})\Big{|}\leq\epsilon\quad\forall s\in S.$ (4)
Equation 4 (see proof in Appendix A.3) shows that for $t\geq\kappa$, acting greedily with respect to $Q_{i}^{*}$ leads to Bayes-optimal behavior, and explicitly maintaining the Bayes-optimal belief over tasks is no longer required. This implies that the experience history $\Upsilon$ can be ignored at that point.
Moreover, Equation 4 implies that for $t<\kappa$,
where $\varepsilon_{i}(\Upsilon)$ denotes the error in $Q$-value estimates as a function of the experience history $\Upsilon$. While this error diminishes as $t\to\infty$, in the short run, a function $f(\Upsilon)$ could be learned to either estimate the error or directly approximate $\bar{V}^{*}(\bar{b})$. The better performance of RL3 could be explained by either the error $\varepsilon_{i}(\Upsilon)$ being simpler to estimate, or the meta-agent’s behavior being more robust to errors in $\varepsilon_{i}(\Upsilon)$ when $Q$-estimates are supplied directly as inputs, compared to errors arising from approximating $\bar{V}^{*}(\bar{b})$ without such inputs. Additionally, this approach benefits from the fact that the convergence rate of $Q$-estimates suggests a natural, predictable schedule of shifting reliance from $f(\Upsilon)$ to $Q_{i}^{t}(s)$ as $t\to\infty$. However, we do not impose this structure explicitly in the network and instead allow it to learn implicitly how much to rely on the $Q$-estimates, informed by the context and supported by the provided action counts as supplementary signals.
Finally, we note that achieving accurate approximation of $\bar{V}^{*}(\bar{b})$ as $t\to\infty$ reduces the error in meta-value estimation for all preceding belief states, since meta-values for consecutive beliefs $\bar{b}$ and $\bar{b}^{\prime}$ are linked via the Bellman equation for BAMDPs (see Appendix A.3):
$\bar{V}^{*}(\bar{b})=\max_{a\in A}\Big{[}\sum_{M_{i}\in\mathcal{M}}\bar{b}(i)R_{i}(s,a)+\gamma\sum_{\bar{\omega}\in\bar{\Omega}}\Pr(\bar{\omega}|\bar{b},a)\bar{V}^{*}(\bar{b}^{\prime})\Big{]},$ (6)
This dependency helps meta-training in RL3 when using temporal-difference learning algorithms. Without conditioning on $Q$-estimates, the error in $\bar{V}^{*}(\bar{b})$ could instead grow as $t\to\infty$, since the meta-critic must condition on an increasingly large experience history $\Upsilon$. This could destabilize meta-value learning across all preceding belief states during meta-training.
4.2 Implementation
@@ 5 Experiments
We compare RL3 to our enhanced implementation of RL2. In our implementation, we replace LSTMs with transformers in both the meta-actor and meta-critic for the purpose of mapping experiences to actions and meta-values, respectively. This is done to improve RL2 ’s ability to handle long-term dependencies instead of suffering from vanishing gradients. Moreover, RL2-transformer trains significantly faster than RL2-LSTM. Second, we include in the state space the total number of interaction steps and the total number of steps within each episode during a meta-episode (see Fig. 1). Third, we use PPO (Schulman et al., 2017) for training the meta actor-critic, instead of TRPO (Schulman et al., 2015). These modifications and other minor-implementation details incorporate the recommendations made by Ni et al. (2022), who show that model-free recurrent RL is competitive with strong Meta-RL approaches such as VariBAD (Zintgraf et al., 2020), if implemented properly. RL3 simply applies the modified version of RL2 to the distribution of value-augmented MDPs explained in section 4.2. Within each VAMDP, our choice of object-level RL is a model-based algorithm to maximize data efficiency – we estimate a tabular model of the environment and run finite-horizon value-iteration using the model. Once again, we emphasize that the core of our approach, which is augmenting MDP states with action-value estimates, is not inherently tied to RL2 and is orthogonal to most other Meta-RL research. VAMDPs can be plugged into any base Meta-RL algorithm with a reasonable expectation of improving it.
@@
Bernoulli Bandits: We use the same setup described by Duan et al. (2016) with $k=5$ arms. To test OOD generalization, we generate bandit tasks by sampling success probabilities from $\mathcal{N}(0.5,0.5)$. We should note that this is an easy domain and serves as a sanity check to ensure that $Q$-value estimates do not hurt RL3, causing inferior performance.
Random MDPs: We use the same setup described by Duan et al. (2016). The MDPs have 10 states, 5 actions, and task horizon 10. The rewards and transition probabilities are drawn from a normal and a flat Dirichlet distribution ($\alpha=1.0$), respectively. OOD test MDPs use Dirichlet $\alpha=0.25$. We should note that this domain is particularly challenging for RL3 due to the high degree of stochasticity and thus the slower convergence rate of $Q$-estimates.
@@ 6 Results
In summary, we observe that beyond matching or exceeding the performance of RL2-transformer in all test domains i) RL3 shows better OOD generalization, which we attribute to the increased generality of the $Q$-value representation, ii) the advantages of RL3 increase with longer interactions periods and less stochastic tasks, which we attribute to the increased accuracy of the $Q$-value estimates, iii) RL3 performs well even with coarse-grained object-level RL over abstract states with substantial computational savings, showing minimal drop in performance in most cases, and iv) RL3 shows faster meta-training.
Figure 3: Results for the MDPs and GridWorlds domains. Figure 3a shows the average cumulative reward (negligible standard error) earned as a fraction of the oracle policy for in-distribution (solid) and OOD (dashed) tasks; Figure 3b shows the fraction of RL2-transformer meta-training iterations that RL3 requires (variance is insignificant across seeds) to match RL2-transformer performance or fully converge, both as functions of the adaptation period $H$. Note the log horizontal axis on both plots. Figure 3c shows the average cumulative reward ($\pm$ standard error) earned by RL2-transformer, RL3 and RL3-coarse agents on several variations of the GridWorlds domain.
@@
MDPs: Figures 3(a) and 3(b) show the results for the MDPs domain. In Figure 3(a), we see that for relatively short budgets, $H\leq 500$, both RL3 and RL2-transformer perform comparably on in-distribution problems, with RL3 performing slightly better on OOD tasks. We suspect that, due to the short budgets and highly stochastic domain, $Q$-estimates do not converge enough to be very useful for RL3 . However, as the budget increases, we see that RL3 continues to improve while RL2-transformer actually becomes worse and the performance gap on both in-distribution and OOD tasks becomes significant. Overall, we see that RL3 preserves asymptotic scaling properties of traditional RL while simultaneously maintaining strong OOD performance. Moreover, RL3 it is able to learn meta-policies much more efficiently. Figure 3(b) shows the number of iterations of PPO RL3 takes to converge completely, as well as to match the performance of RL2-transformer, measured as a fraction of the time it takes for RL2-transformer to converge. This advantage of RL3 is again most pronounced for longer adaptation periods, but we still do observe significant meta-training speedup on even moderate ones. The training curves are available in Appendix C. Overall, it is clear that as adaptation periods grow, RL3 achieves nearer-to-optimal policies in a fraction of the meta-training time and maintains better OOD generalization.
[…]
@@
Fig. 4 shows a sequence of snapshots of a meta-episode where the trained RL3 agent is interacting with an instance of a 13x13 grid. The first snapshot shows the agent just before reaching the goal for the first time. Prior to the first snapshot, the agent had explored many locations in the grid. The second snapshot shows the next episode just after the agent finds the goal, resulting in value estimates being updated using object-level RL for all visited states. Snapshot 3 shows the agent consequently using the $Q$-estimates to navigate to the goal presumably by choosing high-value actions. The agent also explores several new nearby states for which it does not have $Q$-estimates. Snapshot 4 shows the final $Q$-value estimates. A set of short videos of the GridWorlds environment, showing both RL2-transformer and RL3 agents solving the same set of problem instances, can be found at https://youtu.be/eLA_S1BQUYM.
Computation Overhead Considerations: As mentioned earlier, for implementing object-level RL, we use model estimation followed by finite-horizon value-iteration to obtain $Q$-estimates. The computation overhead is negligible for Bandits (5 actions, task horizon = 1) and very little for the MDPs domain (10 states, 5 actions, task horizon 10). For 13x13 GridWorlds (up to 169 states, 5 actions, task horizon = 350), RL3 takes approximately twice the computation time of RL2-transformer per meta-episode. However, RL3-coarse requires only 10% overhead while still outperforming RL2-transformer and retaining more than 90% of the performance of RL3. This demonstrates the utility of state abstractions in RL3 for scaling. Finally, the meta-training sample efficiency demonstrated by RL3 translates directly to wall-time efficiency as training is dominated by gradient computation, not value iteration during data collection in PPO. Our implementation is available at https://github.com/bhatiaabhinav/RL3.
@@ 7 Limitations and Conclusion
To conclude, in this paper, we introduced RL3, a principled hybrid approach that combines the strengths of traditional RL and Meta-RL and provides a more robust and efficient Meta-RL algorithm. We advanced intuitive and theoretical arguments regarding its suitability for Meta-RL and presented empirical evidence to validate those ideas. Specifically, we demonstrated that RL3 holds potential to enhance long-term performance, generalization on out-of-distribution tasks and reducing meta-training time.
@@ Acknowledgments
Supplementary Materials
The following content was not necessarily subject to peer review.
Appendix A Proofs
@@ A.1 Bayes Optimality of $Q$-value Estimates in Bernoulli Multi-armed Bandits
Let the number of times arm $k$ is pulled up to time $T$ be $N_{ik}^{T}$, and the number of successes associated with pulling arm $k$ up to time $T$ be $q_{ik}^{T}$. Given that this is an MDP with just a single state and task horizon of 1, the $Q$-estimate associated with arm $k$ is just the average reward for that action, which is the ratio of successes to counts associated with that action i.e., $Q_{ik}^{T}=\frac{q_{ik}^{T}}{N_{ik}^{T}}$. To reduce the clutter in the notation, we will drop the superscript $T$ for the rest of the subsection. Now,